Good evening. i would like to ask you to help me with this problem
i cannot work out, please would you help me?
(1) John makes business trip from his house to Portland in 2 hours. Three hours later, he returns home in traffic at 20mph less than his speed going. If John is gone a total of 8 hours, how fast did he travel to Portland.

Answer by checkley71(8403)

DISTANCE=RATE*TIME & T=D/R OR 2=D/R OR D=2R GOING TO PORTLAND & (8-2-3)=D/(R-20) OR 3=D/(R-20) OR D=3R-60 RETURNING FROM PORTLAND THUS 2R=3R-60 OR R=60 MPH
ON THE TRIP TO PORTLAND